OpenStudy (anonymous):

help! what is -5/8w=-9

OpenStudy (anonymous):

5/72

OpenStudy (vijay):

w =5/72

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

-5/8w=-9, now bring -9 on the other side.when this happens, -9 goes in the denominator.so we get -5/(-9*8w)=1 now we need to isolate w, so take it on the other sie, when it goes on the other side, it is multiplied, so we get -5/(-9*8)=w w=-5/-72 = 5/72

OpenStudy (anonymous):

\[\text{Solve}\left[-\frac{5}{8}w\text{ = }-9,w\right]=\frac{72}{5}\]

OpenStudy (anonymous):

wait....is it 72/5 or 5/72???

OpenStudy (anonymous):

\[-\frac{5}{8}*\frac{72}{5}=-\frac{72}{8}=-9 \]